Natalie's father is 4 times as old as Natalie. After 5 years, the father will be three times as old as Natalie. Find their present ages.
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Natalie's father is 4 times as old as Natalie. After 5 years, the father will be three times as old as Natalie. Find their present ages.
------------------
F = 4N
F+5 = 3*(N+5)

Answer by ikleyn(52787)

.

If Natalie's age is x years, then the father's age is 4x, according to the condition.


After 5 years


    3*(x+5) = 4x + 5

    3x + 15 = 4x + 5

    15 - 5 = 4x - 3x

    10     = x.


ANSWER.  Natalie is 10 years old.  Her father is 4x = 4*10 = 40 years old.

Solved.

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.